A savings plan is set up so that $100 is placed into an account on the 1st and 15th of each month. The first payment was made on January 1, 2010 and this continued until December 1, 2013. Then no money was deposited for the rest of December 2013 until March 1, 2014 when the $100 payments were continued until the last payment was made on February 1, 2017. Interest is 2.5% effective annual interest. a. Draw a careful number line and indicate the number of payments that were made before the missed payment, the number of missed payments and the number of payments after the missed payments. Remember that the payments are twice per month. b. Find the annual rate compounded twice per month. c. Find the amount in the account after the payment is made on December 1, 2013. d. Find the amount in the account after the payment is made on March 15, 2014. e. Find the amount in the account on February 15, 2017
Answer #1

Part (a)

Timeline and payment

Payment no. Date Payment
1 1-Jan-10 100
2 15-Jan-10 100
3 1-Feb-10 100
4 15-Feb-10 100
5 1-Mar-10 100
6 15-Mar-10 100
7 1-Apr-10 100
8 15-Apr-10 100
9 1-May-10 100
10 15-May-10 100
11 1-Jun-10 100
12 15-Jun-10 100
13 1-Jul-10 100
14 15-Jul-10 100
15 1-Aug-10 100
16 15-Aug-10 100
17 1-Sep-10 100
18 15-Sep-10 100
19 1-Oct-10 100
20 15-Oct-10 100
21 1-Nov-10 100
22 15-Nov-10 100
23 1-Dec-10 100
24 15-Dec-10 100
25 1-Jan-11 100
26 15-Jan-11 100
27 1-Feb-11 100
28 15-Feb-11 100
29 1-Mar-11 100
30 15-Mar-11 100
31 1-Apr-11 100
32 15-Apr-11 100
33 1-May-11 100
34 15-May-11 100
35 1-Jun-11 100
36 15-Jun-11 100
37 1-Jul-11 100
38 15-Jul-11 100
39 1-Aug-11 100
40 15-Aug-11 100
41 1-Sep-11 100
42 15-Sep-11 100
43 1-Oct-11 100
44 15-Oct-11 100
45 1-Nov-11 100
46 15-Nov-11 100
47 1-Dec-11 100
48 15-Dec-11 100
49 1-Jan-12 100
50 15-Jan-12 100
51 1-Feb-12 100
52 15-Feb-12 100
53 1-Mar-12 100
54 15-Mar-12 100
55 1-Apr-12 100
56 15-Apr-12 100
57 1-May-12 100
58 15-May-12 100
59 1-Jun-12 100
60 15-Jun-12 100
61 1-Jul-12 100
62 15-Jul-12 100
63 1-Aug-12 100
64 15-Aug-12 100
65 1-Sep-12 100
66 15-Sep-12 100
67 1-Oct-12 100
68 15-Oct-12 100
69 1-Nov-12 100
70 15-Nov-12 100
71 1-Dec-12 100
72 15-Dec-12 100
73 1-Jan-13 100
74 15-Jan-13 100
75 1-Feb-13 100
76 15-Feb-13 100
77 1-Mar-13 100
78 15-Mar-13 100
79 1-Apr-13 100
80 15-Apr-13 100
81 1-May-13 100
82 15-May-13 100
83 1-Jun-13 100
84 15-Jun-13 100
85 1-Jul-13 100
86 15-Jul-13 100
87 1-Aug-13 100
88 15-Aug-13 100
89 1-Sep-13 100
90 15-Sep-13 100
91 1-Oct-13 100
92 15-Oct-13 100
93 1-Nov-13 100
94 15-Nov-13 100
95 1-Dec-13 100
96 15-Dec-13
97 1-Jan-14
98 15-Jan-14
99 1-Feb-14
100 15-Feb-14
101 1-Mar-14 100
102 15-Mar-14 100
103 1-Apr-14 100
104 15-Apr-14 100
105 1-May-14 100
106 15-May-14 100
107 1-Jun-14 100
108 15-Jun-14 100
109 1-Jul-14 100
110 15-Jul-14 100
111 1-Aug-14 100
112 15-Aug-14 100
113 1-Sep-14 100
114 15-Sep-14 100
115 1-Oct-14 100
116 15-Oct-14 100
117 1-Nov-14 100
118 15-Nov-14 100
119 1-Dec-14 100
120 15-Dec-14 100
121 1-Jan-15 100
122 15-Jan-15 100
123 1-Feb-15 100
124 15-Feb-15 100
125 1-Mar-15 100
126 15-Mar-15 100
127 1-Apr-15 100
128 15-Apr-15 100
129 1-May-15 100
130 15-May-15 100
131 1-Jun-15 100
132 15-Jun-15 100
133 1-Jul-15 100
134 15-Jul-15 100
135 1-Aug-15 100
136 15-Aug-15 100
137 1-Sep-15 100
138 15-Sep-15 100
139 1-Oct-15 100
140 15-Oct-15 100
141 1-Nov-15 100
142 15-Nov-15 100
143 1-Dec-15 100
144 15-Dec-15 100
145 1-Jan-16 100
146 15-Jan-16 100
147 1-Feb-16 100

the number of payments that were made before the missed payment = 95

the number of missed payments = 5

and the number of payments after the missed payments = 47

Part (b)

Let r be the annual rate compounded twice per month

EAR = 2.5% = (1 + r / 24)24 - 1

Hence, r = [(1 + 2.5%)1/24 - 1] x 24 = 0.1029%

Part (c)

the amount in the account after the payment is made on December 1, 2013 = FV (Rate, Nper, PMT, PV) = FV (0.1029%, 94, -100, -100) = $9,974.64

Part (d)

the amount in the account after the payment is made on March 1, 2014 = $9,974.64 x (1 + 0.1029%)6 + 100 = $10,136.41

Part (e)

the amount in the account on the last payment date of February 1, 2016 = FV (Rate, Nper, PMT, PV) = FV (0.1029%, 46, -100, -10136.41) = $15,335.84